the growth of a town's population is modeled by p=104e^0.03t how large will the population be when t=24
Answer by checkley77(12844)

p=104e^0.03t
p=104*2.7182^(.03*24)
p=104*2.7182^.72
p=104*1.71
p=178 ans.